a train starts with some passengers. at the first stop one third of the passengers got down and 96 more passengers boarded the train. If there are 256 passengers present in the train, the number of passengers in the beginning was
1
Expert's answer
2012-11-08T09:56:30-0500
Let the amount of passengers at start of train X
At the station 1/3& (1/3X)of passengers came out and 96 came in the amount of passengers is:
X-1/3X+96 and equal to 256
X-1/3X+96=256
2/3X=256-96
2/3X=160
2X=480
X=240
Answer: at the start there was 240 passengers in train